# Hyperbolic embedding of multilayer networks

### v1

@@ II.5 Layers extraction and radii attribution
Once the non-linear dimensionality reduction has been performed, nodes retain their angular coordinates in the two-dimensional embedding, while their radial positions are reassigned based on their centrality within the network. As illustrated in Fig. 1, the initial step involves extracting the positions of the nodes for each layer and setting all radial coordinates to $1$. Radial positions are then reassigned according to each node’s centrality, measured by its degree. To obtain a more informative estimate of centrality, we use the weighted degree $w_{i}$, which accounts for the strength of connections rather than merely their number. Notably, this centrality measure reflects any pre-weighting applied to the edges; in the absence of such preprocessing, the original edge weights are used. Although weighted degree is employed here, alternative centrality metrics may also be used, depending on the specific application or network characteristics[25, 32]. Given a node $i$ with (weighted) degree $w_{i}$, its radial coordinate is updated as follows:
@@
where $r_{i}$ denotes the radial coordinate of node $i$, $\tanh$ is the hyperbolic tangent function, and $\beta$ is a scaling parameter. This formulation ensures that all nodes are embedded within the Poincaré disk, with nodes of degree zero mapped at the boundary ($r=1$) and the most central nodes positioned closer to the origin. The parameter $\beta$ controls the rate at which the radius decreases with increasing centrality and plays a role analogous to the popularity fading parameter in the original coalescent embedding method.
@@ III.1 Multilayer community structure – Multilayer SBM example
$p_{i,j}=\begin{cases}p_{\neq}&\text{if }c_{i}\neq c_{j}\text{ and }l_{i}=l_{j},\\ p_{=}&\text{if }c_{i}=c_{j}\text{ and }l_{i}=l_{j},\\ \frac{p_{\neq}}{\alpha}&\text{if }c_{i}\neq c_{j}\text{ and }l_{i}\neq l_{j},\\ \frac{p_{=}}{\alpha}&\text{if }c_{i}=c_{j}\text{ and }l_{i}\neq l_{j}.\\ \end{cases}$ (15)
@@
To demonstrate the effectiveness of our approach in revealing community structure, we compared the results of the hyperbolic embedding in the Poincaré disk under two settings: i) independent embedding of each layer without cross-layer information, and ii) joint embedding of all layers using our multilayer method. Remarkably, even under weak inter-layer coupling ($\alpha=10$), the multilayer approach reveals well-separated community clusters, unlike the independent embeddings where the structure appears more dispersed. These results, shown in Fig. 2, highlight the advantage of the multilayer embedding in networks with latent community structure. When using our multilayer framework, the communities are clearly delineated in the hyperbolic space, whereas the independent approach requires stronger intra-community and lower inter-community connectivity to achieve comparable separation.
@@ III.2 Nodes localization – An example on brain epilepsy networks
To illustrate the practical utility of our approach, we computed the hyperbolic Gaussian distribution of node positions (specifically from the left temporal lobes) within the Poincaré disk (See Sec. II.6). We evaluated two embedding strategies: our proposed multilayer method, in which each layer represents a different patient, and an independent embedding approach, wherein each patient is embedded separately and the resulting configurations are subsequently aligned via rotational transformation.
As shown in Fig. 3, the multilayer embedding produces a more consistent and spatially coherent localization of the temporal lobe regions compared to the independent embedding strategy. From this result we observe that, compared with independently embedding each layer, the nodes within the left temporal lobe are more tightly clustered under our method, yielding a sharper Gaussian distribution. This enhanced localization highlights the capacity of the multilayer embedding to integrate information across patients, thereby generating more stable and reliable individual embeddings. Moreover, the method clearly uncovers distributional differences between patients with left temporal epilepsy and healthy controls, underscoring its potential utility in clinical settings.
Figure 3: Hyperbolic embedding of patients with left temporal lobe epilepsy (red points) and controls (blue points). Visualization of the Gaussian distributions of node positions corresponding to the left temporal lobes within the Poincaré disk, using two embedding strategies: A) the proposed multilayer embedding approach, where each layer represents an individual patient; and B) an independent embedding of each patient followed by post hoc rotational alignment.
@@ III.3 Impact of the $\mu$ coupling parameter
To assess the impact of the coupling parameter $\mu$, we performed the embedding over a range of $\mu$ values for a two-layer graph (a multilayer SBM model), computing the corresponding $G_{\text{score}}$ in each case. Interestingly, we observed a sharp transition in the behavior of $G_{\text{score}}$ as $\mu$ increases. Beyond a critical value, denoted $\mu^{*}$, the score drops significantly and stabilizes at a plateau value, referred to as ${G_{\text{score}}}^{*}$ (see Fig. 4). Plateau is stable even when $\mu$ takes large values such as $\mu\to 10^{4}\mu^{*}$ (result not shown).
@@
The global plateau value ${G_{\text{score}}}^{*}$ also appears to carry physical significance. While the critical coupling $\mu^{*}$ is governed by the average edge weight, the value of ${G_{\text{score}}}^{*}$ seems to reflect the similarity between the connectivity distributions of the layers. To assess this hypothesis, we perturbed the connection probabilities in the second layer (i.e., modifying $p_{=}$ and $p_{\neq}$) while keeping the first layer unchanged. As shown in Fig. 4(b), the relationship between $\mu^{*}$ and the mean edge weight remains robust to these perturbations, indicating that the phase transition is not disrupted. However, the plateau value ${G_{\text{score}}}^{*}$ increases when the two layers become less similar, compared to the case with identical probability distributions (Fig. 4(a)). These results suggest that ${G_{\text{score}}}^{*}$ can serve as a quantitative indicator of interlayer similarity: the more similar the layers, the lower the score.
@@ IV Discussion
Our method is entirely data-driven, in contrast to model-based approaches such as Mercator [23], which rely on generative assumptions. This distinction renders our framework more flexible and widely applicable, particularly in empirical settings where model parameters may be unknown or difficult to estimate.
@@
Another central contribution of this work is the analysis of the coupling parameter $\mu$, which modulates the influence of inter-layer interactions in the global embedding. This parameter governs the influence of inter-layer interactions in the global embedding. We show that the alignment quality of the embeddings undergoes a clear transition as $\mu$ increases: below a critical threshold $\mu^{*}$, embeddings remain poorly aligned, whereas above $\mu^{*}$, a plateau is reached, indicating successful integration of layers.. Notably, we observe that $\mu^{*}$ scales with the average edge weight of the network, suggesting that optimal coupling reflects the intrinsic connectivity strength within the layers. Furthermore, the plateau value of the alignment score serves as a proxy for inter-layer similarity, with lower scores corresponding to more structurally homogeneous layers. These findings suggest that $\mu$ acts not only as a tunable parameter for controlling multilayer integration,but also as a diagnostic tool for uncovering latent geometric and structural properties of multilayer systems.

### v2

@@ II.5 Layers extraction and radii attribution
Once the non-linear dimensionality reduction has been performed, nodes retain their angular coordinates in the two-dimensional embedding, while their radial positions are reassigned based on their centrality within the network. As illustrated in Fig. 1, the initial step involves extracting the positions of the nodes for each layer and setting all radial coordinates to $1$. Radial positions are then reassigned according to each node’s centrality, measured by its degree. To obtain a more informative estimate of centrality, we use the weighted degree $w_{i}$, which accounts for the strength of connections rather than merely their number.The weighted degree of a node is defined as the sum of the weights of its incident edges, providing a measure of the node’s overall connectivity strength within the weighted network. Notably, this centrality measure reflects any pre-weighting applied to the edges; in the absence of such preprocessing, the original edge weights are used. Although weighted degree is employed here, alternative centrality metrics may also be used, depending on the specific application or network characteristics[25, 32]. Given a node $i$ with (weighted) degree $w_{i}$, its radial coordinate is updated as follows:
@@
where $r_{i}$ denotes the radial coordinate of node $i$, $\tanh$ is the hyperbolic tangent function, and $\beta$ is a scaling parameter. This formulation ensures that all nodes are embedded within the Poincaré disk, with nodes of degree zero mapped at the boundary ($r=1$) and the most central nodes positioned closer to the origin. The parameter $\beta$ controls the rate at which the radius decreases with increasing centrality and plays a role analogous to the popularity fading parameter in the original coalescent embedding method. A practical strategy for setting $\beta$ consists in prescribing the radius $r_{i}$ of a node $i$ with maximal (weighted) degree $w_{i}$, such that it lies within a desired distance from the origin (for instance, $r_{i}=0.1/0.2$) and solving $\beta=w_{i}/\text{arctanh}(r_{i})$. This ensures that node radii span a broad portion of the disk while maintaining an interpretable mapping between centrality and geometry.
@@ III.1 Multilayer community structure – Multilayer SBM example
[…]
@@
To demonstrate the effectiveness of our approach in revealing community structure, we compared the results of the hyperbolic embedding in the Poincaré disk under two settings: i) independent embedding of each layer without cross-layer information, and ii) joint embedding of all layers using our multilayer method. Remarkably, even under weak inter-layer coupling ($\alpha=10$), the multilayer approach reveals well-separated community clusters, unlike the independent embeddings where the structure appears more dispersed. These results, shown in Fig. 2, highlight the advantage of the multilayer embedding in networks with latent community structure. When using our multilayer framework, the communities are clearly delineated in the hyperbolic space, whereas the independent approach requires stronger intra-community and lower inter-community connectivity to achieve comparable separation. This is due to the fact that a lower inter-community connectivity enhances the contrast between intra- and inter-community edge probabilities, which is essential for the independent embedding to produce a clear separation of communities. When this contrast diminishes and the connection probabilities become similar, the network approaches a regime akin to an Erdős–Rényi graph, where communities are less distinguishable without supplementary interlayer information. When replacing Isomap with Laplacian eigenmaps on the same dataset, the multilayer embedding still preserves community separation, while the performance of the independent embedding further deteriorates with the spectral method.
@@
Other multilayer/multiplex network models exist, including the Geometric Multiplex Model (GMM) [35, 29]. We applied our embedding method to networks generated using the GMM model and compared the pairwise hyperbolic distances between nodes in the original configuration and the embedded version. Using a 4-layer network with 100 nodes, we observed a mean Spearman correlation of 0.79 between the original and embedded hyperbolic distances, indicating that the embedding effectively captures the network geometry of the network.”
@@ III.2 Nodes localization – An example on brain epilepsy networks
To illustrate the practical utility of our approach, we computed the hyperbolic Gaussian distribution of node positions (specifically from the left temporal lobes) within the Poincaré disk (See Sec. II.6). We evaluated two embedding strategies: our proposed multilayer method, in which each layer represents a different patient, and an independent embedding approach, wherein each patient is embedded separately and the resulting configurations are subsequently aligned via rotational transformation. The alignment is performed by rotating each embedding (beyond the reference layer) to minimize the sum of hyperbolic distances between corresponding nodes, effectively minimizing the $G_{\text{score}}$ defined Eq. 16. This procedure, similar to the one described in [37], ensures comparability across layers while preserving the geometric invariance of the embedding space.
As shown in Fig. 3, the multilayer embedding produces a more consistent and spatially coherent localization of the temporal lobe regions compared to the independent embedding strategy. From this result we observe that, compared with independently embedding each layer, the nodes within the left temporal lobe are more tightly clustered under our method, yielding a sharper Gaussian distribution. This enhanced localization highlights the capacity of the multilayer embedding to integrate information across patients, thereby generating more stable and reliable individual embeddings. Moreover, the method clearly uncovers distributional differences between patients with left temporal epilepsy and healthy controls, underscoring its potential utility in clinical settings. Permutation tests, where group labels were randomly reassigned, resulted in entropy distributions for both groups converging towards a common intermediate value, distinct from the lower entropy observed in patients (a more concentrated node positions) and higher entropy in controls (more dispersed configurations). This indicates that the tighter clustering in patients is not an artifact of the embedding method but reflects genuine group-specific network organization.
Figure 3: Hyperbolic embedding of patients with left temporal lobe epilepsy (red points) and controls (blue points). Visualization of the Gaussian distributions of node positions corresponding to the left temporal lobes (2 nodes per patients) within the Poincaré disk, using two embedding strategies: A) the proposed multilayer embedding approach, where each layer represents an individual patient; and B) an independent embedding of each patient followed by post hoc rotational alignment.
@@ III.3 Impact of the $\mu$ coupling parameter
To assess the impact of the coupling parameter $\mu$, we performed the embedding over a range of $\mu$ values for a two-layer graph (a multilayer SBM model), computing the corresponding $G_{\text{score}}$ in each case. Interestingly, we observed a sharp transition in the behavior of $G_{\text{score}}$ as $\mu$ increases. Beyond a critical value, denoted $\mu^{*}$, the score drops significantly and stabilizes at a plateau value, referred to as $G_{\text{score}}$$^{*}$ (see Fig. 4). Plateau is stable even when $\mu$ takes large values such as $\mu\to 10^{4}\mu^{*}$ (result not shown). This indicates that, to reach a stable regime on the plateau, the parameter $\mu$ must satisfy $\mu\gg\mu^{*}\simeq$ the mean edge weight on the graph. Since the alignment remains stable once the plateau is reached, choosing a sufficiently high value of $\mu$ ensures convergence without adverse effects.
@@
The global plateau value $G_{\text{score}}$$^{*}$ also appears to carry physical significance. While the critical coupling $\mu^{*}$ is governed by the average edge weight, the value of $G_{\text{score}}$$^{*}$ seems to reflect the similarity between the connectivity distributions of the layers. To assess this hypothesis, we perturbed the connection probabilities in the second layer (i.e., modifying $p_{=}$ and $p_{\neq}$) while keeping the first layer unchanged. As shown in Fig. 4(b), the relationship between $\mu^{*}$ and the mean edge weight remains robust to these perturbations, indicating that the phase transition is not disrupted. However, the plateau value $G_{\text{score}}$$^{*}$ increases when the two layers become less similar, compared to the case with identical probability distributions (Fig. 4(a)). These results suggest that $G_{\text{score}}$$^{*}$ can serve as a quantitative indicator of interlayer similarity: the more similar the layers, the lower the score. To further validate this interpretation, we compared $G_{\text{score}}$$^{*}$ with the graph edit distance [38] in a controlled perturbation experiment, where edges were randomly added and removed from an initial network. The two metrics showed a strong linear relationship, with a Pearson correlation coefficient of 0.969, confirming that $G_{\text{score}}$$^{*}$ reliably captures structural dissimilarities between layers.
@@ IV Discussion
Our method is entirely data-driven and does not rely on explicit generative assumptions, in contrast to model-based approaches such as Mercator [23], which infer latent geometric parameters through a probabilistic framework grounded in network generative models. While Mercator provides robust parameter estimation and interpretable embeddings even in weakly geometric regimes, our approach offers complementary advantages by allowing flexible and direct embedding without requiring prior model parameter estimation. This makes our framework particularly suitable for empirical settings where model assumptions may not hold or where parameters are challenging to infer, such as in heterogeneous or partially observed multilayer networks.
@@
Another central contribution of this work is the analysis of the coupling parameter $\mu$, which modulates the influence of inter-layer interactions in the global embedding. This parameter governs the influence of inter-layer interactions in the global embedding. We show that the alignment quality of the embeddings undergoes a clear transition as $\mu$ increases: below a critical threshold $\mu^{*}$, embeddings remain poorly aligned, whereas above $\mu^{*}$, a plateau is reached, indicating successful integration of layers.. Notably, we observe that $\mu^{*}$ scales with the average edge weight of the network, suggesting that optimal coupling reflects the intrinsic connectivity strength within the layers. Furthermore, the plateau value of the alignment score $G_{\mathrm{score}}$ serves as a proxy for inter-layer similarity, with lower scores corresponding to more structurally homogeneous layers. These findings suggest that the coupling parameter $\mu$ acts not only as a tunable parameter for controlling multilayer integration, but also as a diagnostic indicator of the underlying network structure. Specifically, the optimal value $\mu$ must be sufficiently large relative to the mean edge weight of the network to ensure stable alignment across layers. Furthermore, the plateau value of the alignment score, $G_{\text{score}}^{*}$, quantitatively reflects the degree of similarity or dissimilarity between layers, thus providing insight into the latent geometric and structural organization of the multilayer system.
@@
An important limitation of the proposed method concerns its computational scalability. The approach requires constructing and processing a global connectivity matrix of size $M\times M$, where $M=N\times L$, with $N$ the number of nodes per layer and $L$ the number of layers. Consequently, the computational complexity scales approximately as $\mathcal{O}(M^{2})$, which can quickly become prohibitive for multilayer networks with a large number of layers or nodes. This limitation restricts the applicability of our method to moderately sized multilayer systems, particularly in the context of temporal networks with high temporal resolution, where the number of layers (snapshots) is very large. Addressing this challenge will require algorithmic optimizations or approximation strategies, representing a promising direction for future work.
@@ Acknowledgements.
We thank Marián Ángeles Serrano for kindly sharing the implementation of the Geometric Multiplex Model (GMM), which was used in our simulations.